A long, thin solenoid has 930 turns per meter and radius 2.90 cm . The current in the solenoid is increasing at a uniform rate of 63.0 A/s .

Part A

What is the magnitude of the induced electric field at a point 0.520 cm from the axis of the solenoid?

Part B

What is the magnitude of the induced electric field at a point 1.10 cm from the axis of the solenoid?
Answer #1

electric field is given by

E = 0.5*uo n r (dI/dt)

a) if r = 0.52* 10^-2 m

putting in the above equation

E = 0.5* 4 * 3.14* 10^-7* 930* 0.52* 10^-2* 63

E = 1.91* 10^-4 N/C

========

b)

similarly put, r = 1.1* 10^-2 m

E = 4.05* 10^-4 N/C

========

Goodluck